# Differentiable Discrete Elastic Rods for Real-Time Modeling of Deformable Linear Objects

### v1

@@ A.2 Summary and Discussion of Improved Inextensiblity Enforcement with PBD
Algorithm 2 summarizes our proposed method to enforce inextensibility while preserving momentum. Note that in practice, the while loop in Algorithm 2 typically converges after two iterations for $\epsilon=0.05$. Once we have the output from Algorithm 2, we update the velocities of the vertices to reflect the new vertex locations (i.e., Line 5 of Algorithm 1). Further, as shown in Table 6, skipping Algorithm 2 in Algorithm 1 and relying solely on DNNs to capture the inextensibility of DLOs can lead to simulation instability. This instability arises because modeling stiff behavior, such as the inextensibility of DLOs, makes the learning process highly sensitive to inputs and hinders effective gradient propagation [8, 11]. On the contrary, our approach explicitly imposes inextensibility, and seperates DNNs from the complications associated with learning stiff behavior.
A.3 Training Setup
@@
Appendix B Experiemental Details
B.1 Hardware Parameters
@@ B.3 Ablation Study: DLO4 and DLO5
B.4 Ablation Study: Inextensibility Enforcement with Learning and Single-step Training
@@
B.5 ARMOUR
To perform a shape matching task, we rely upon ARMOUR[41], an optimization-based motion planning and control framework. The goal of the shape matching task is to use a robot arm to manipulate the DLO from an initial configuration to a predefined target configuration. To accomplish this goal, ARMOUR performs planning in a receding horizon fashion. During each planning iteration, ARMOUR selects a trajectory to follow by solving an optimization problem. More details about ARMOUR’s trajectory parameterization and associated closed loop controller can be found in [41, Section IX]. The cost function minimizes the distance between the predicted DLO configuration given a chosen robot trajectory and a target DLO configuration. The predicted state of the DLO is computed via a DLO modeling technique.
Appendix C DLO Tracking with Modeling
This section first discusses the difficulties of incorporating existing framework with modeling for long-time DLO tracking under occlusion. It then proposes a novel perception pipeline, which is utilized in Section 5.2.
C.1 DLO Tracking Review
@@
C.2 DLO Tracking with DEFORM
@@
C.3 State Estimation in the Presence of Occlusions
@@
C.4 Initial State Estimation under Occlusion

### v3

@@ B.3 Summary and Discussion of Improved Inextensiblity Enforcement with PBD
Algorithm 2 summarizes our proposed method to enforce inextensibility while preserving momentum. Note that in practice, the while loop in Algorithm 2 typically converges after two iterations for $\epsilon=0.05$. Once we have the output from Algorithm 2, we update the velocities of the vertices to reflect the new vertex locations (i.e., Line 5 of Algorithm 1). Further, as shown in Table 6, skipping Algorithm 2 in Algorithm 1 and relying solely on DNNs to capture the inextensibility of DLOs can lead to simulation instability. This instability arises because modeling stiff behavior, such as the inextensibility of DLOs, makes the learning process sensitive to inputs and hinders effective gradient propagation. A key issue with using neural network predictions in position-based dynamics (PBD) is that the dynamics/data lie on a low-dimensional manifold within a high-dimensional space. Over several open-loop predictions, the network’s prediction errors can push the system state away from this manifold, leading to deterioration in the network’s predictions. This occurs because the network has never encountered these “off-manifold” points during training. Algorithm 2 is designed to ensure that model inputs remain on the manifold on which the training data was collected, thereby mitigating these issues.
B.4 Training Setup
@@
B.5 Discussion of Neural Network Utilization in (7)
This section provides an additional discussion of using neural networks in (7) and (8). Incorporating a GNN enhances prediction accuracy by addressing inaccuracies arising from spatial and temporal discretization. This network can effectively identify and adjust for these imperfections, as demonstrated in the ablation study. As the training involves a multi-step process, the GNN allows DEFORM to leverage gradients backpropaged from future prediction losses, which enables better informed and proactive corrections. GNN can implicitly capture variations (e.g., shearing) in residual energy that are not directly modeled by the traditional DER framework.
However, the integration of a neural network also introduces several challenges. Primarily, it increases the computational load. Moreover, the addition of a neural network, particularly in capturing and correcting impulse vectors, introduces complexity in ensuring that the network’s outputs remain physically plausible and consistent with the DER’s modeling output.
B.6 Comparison of DER to other DLO dynamics formulations
Originally derived from computer graphics, the DER approaches uses Lagrangian mechanics to model the dynamics of a chain of point masses, employing a unique choice of coordinates that effectively captures bending and twisting dynamics. This method contrasts with recent studies that apply Newtonian multi-body dynamics using Featherstone’s algorithms [44] and [45]. For instance, [45] models DLO dynamics using a chain of rigid bodies defined in minimal coordinates like angles between body elements, which simplifies the computation by making forward dynamics explicit and eliminating the need for implicit optimization or additional constraint enforcement algorithms. While both Newtonian and Lagrangian dynamics conserve energy, their methods for deriving conservative forces differ, necessitating a deeper discussion on the advantages and limitations of each approach.
From a Lagrangian perspective, DER’s innovative coordinate choice and post-integration constraint enforcement allow for the modeling of complex, stiff behaviors through various optimization solvers, enhancing stability and accuracy crucial for DLO simulations. The accurate and stable model offers a high-fidelity prior for learning residual behavior, significantly enhancing sampling efficiency and performance as shown in the ablation study. However, this can introduce significant computational complexity and burden due to the need to solve implicit optimization problems and enforce constraints. In contrast, Newtonian dynamics offer simpler, faster computations due to their explicit nature and minimal coordinate system, facilitating the integration of complex neural networks. However, this can sometimes result in simulation instabilities, particularly when modeling stiff materials, and may struggle to capture complex deformable behaviors without extensive modifications. The pros and cons of each method illustrate a clear trade-off between computational simplicity and the ability to handle complex behaviors. Overall, there are certainly benefits to using Newtonian multi-body dynamics that warrant further exploration.
Appendix C Experimental Details
C.1 Hardware Parameters
@@ C.3 Ablation Study: DLO4 and DLO5
C.4 Ablation Study: Inextensibility Enforcement with Learning and Single-step Training
@@
C.5 Effect of Number of Vertices: Accuracy and Computational Efficiency
To further investigate the impact of vertices number on accuracy and computational efficiency, we conducted additional experiments using DLO 1. Figure 8 illustrates that increasing the number of vertices enhances prediction accuracy by allowing DEFORM to capture finer details due to improved resolution. However, this increase in detail is at the cost of reduced computational efficiency. In our case we chose the number of vertices to achieve optimal prediction accuracy while ensuring real-time inference capabilities at a 100Hz frequency.
Figure 8: Accuracy and Compuational Time vs. Vertices Number
C.6 ARMOUR
To perform a shape matching task, we rely upon ARMOUR[37], an optimization-based motion planning and control framework. The goal of the shape matching task is to use a robot arm to manipulate the DLO from an initial configuration to a predefined target configuration. To accomplish this goal, ARMOUR performs planning in a receding horizon fashion. During each planning iteration, ARMOUR selects a trajectory to follow by solving an optimization problem. ARMOUR represents joint trajectories of the robot as polynomials, where the coefficients are considered as the decision variables of the optimization. One end of the rope is rigidly attached to the end effector of the robot, whose configuration can be computed by forward kinematics. By plugging the end effector configuration at the end of the trajectories into DEFORM, we are able to get a prediction of the rope configuration at the end of the robot motion. The cost function for the optimization problem to minimize is then defined as the norm distance between the predicted rope configuration and the target rope configuration. Joint position, velocity, and torque limits of the arm are considered as constraints of the optimization problem in ARMOUR’s formulation, which ensures that the robot motion is feasible. The solution to this optimization problem is then the joint trajectories of the robot that manipulate the rope to approach the desired configuration. To command the robot to track the solution trajectories, we use the controller proposed in ARMOUR. This process is conducted repeatedly until the desired configuration is reached, or a planning iteration limit is reached resulting in a failure. More details about ARMOUR’s trajectory parameterization and associated closed loop controller can be found in [37, Section IX]. The cost function minimizes the distance between the predicted DLO configuration given a chosen robot trajectory and a target DLO configuration. The predicted state of the DLO is computed via a DLO modeling technique.
Appendix D DLO Tracking with Modeling
This section first discusses the difficulties of incorporating existing framework with modeling for long-time DLO tracking under occlusion. It then proposes a novel perception pipeline, which is utilized in Section 4.
D.1 DLO Tracking Review
@@
D.2 DLO Tracking with DEFORM
@@
D.3 State Estimation in the Presence of Occlusions
@@
D.4 Initial State Estimation under Occlusion